# Distributionally Robust Markov Games with Average Reward

### v1

@@ Appendix B Proof of Section 4
for some positive distribution matrix $\mu$. Similarly, it implies $x(s)=x(s^{\prime})$, and thus $h^{\pi^{*}}_{\mathsf{P}^{*}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}=ce$. This further implies that all inequalities in (42) are equations, thus
$h^{\pi^{*}}_{\mathsf{P}^{*}}(s)=\max_{a}\left\{r(s,a)-g^{\pi^{*}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}})\right\},$ (45)
hence $(h^{\pi^{*}}_{\mathsf{P}^{*}},g^{\pi^{*}}_{\mathsf{P}^{*}})$ is a solution to (6), completing the proof of the first part. The remaining proof follows in a similar way from Wang et al. (2023b). ∎
Appendix C Proof of Section 5
@@
If for any agent $i$ and any joint policy $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, convex, and $\text{BR}_{i}(\cdot)$ is upper semi-continuous, then BR also satisfies the conditions in Lemma 5.3
@@
(3). Consider two convergent policy sequences $\{\pi^{n}\}\to\pi$ and $\{\mu^{n}\}\to\mu$, with $\mu^{n}\in\text{BR}(\pi^{n})$, we need to show that $\mu\in\text{BR}(\pi)$. Note that $\pi^{n}\to\pi$ implies that $\pi^{n}_{i}\to\pi_{i}$ for any $i$, and similarly $\mu^{n}\to\mu$. Since $\mu_{i}^{n}\in\text{BR}_{i}(\pi_{-i}^{n})$, by the semi-continuity of $\text{BR}_{i}$, we have $\mu_{i}\in\text{BR}_{i}(\pi_{-i})$. Moreover note that
@@ C.1 Convexity
Now from Theorem 4.3, $g=g^{\pi}=g^{*}_{\mathcal{P}}$, thus it holds that
@@
Define $\Pi^{*}\triangleq\{\pi\ |\ \text{optimal policies of }g^{\pi}_{\mathcal{P}}\}$. Then $\Pi^{*}$ is convex.
@@
Let $\pi_{1},\pi_{2}\in\Pi^{*}$ and define $\tilde{\pi}=\alpha\pi_{1}+(1-\alpha)\pi_{2}$ for some $\alpha\in[0,1].$ Then by Theorem 4.3,
@@
Consequently, $(g,h)$ is also a solution to the robust Bellman equation w.r.t. $\tilde{\pi}$. This implies that $g^{\tilde{\pi}}=g=g^{*}$, and thus that $\tilde{\pi}$ is also optimal.
@@ C.2 Semi-continuity
Consider a policy sequence $\{\pi_{n}\}\rightarrow\pi$, and the sequence of induced transition matrices $\{\mathsf{P}_{\pi_{n}}\}$. It holds that $\mathsf{P}_{\pi_{n}}^{*}\rightarrow\mathsf{P}_{\pi}^{*}$ uniformly on $\mathcal{P}$.
@@
See Prop 8.4.6 in Puterman (1994) for the proof. ∎
@@
Theorem C.6.
(Semi-continuity) Let $\pi_{n}\rightarrow\pi$ and $x_{n}\rightarrow x$, where $\pi_{n}:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{-i})$ and $x_{n}:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{i})$. If $x_{n}\in\text{BR}_{i}(\pi_{n}),$ then $x\in\text{BR}_{i}(\pi).$
@@
Moreover, since $h_{n}$ is some relative function Wang et al. (2023a), $h_{n}||_{\infty}\leq H.$ Then, there exists some subsequence $\{n_{m}\},$ s.t. $h_{n_{m}}\rightarrow h$ for some $h$, and $g^{(x_{n},\pi_{n})}_{\mathcal{P}}\rightarrow g$ for some $g$. Then, we can take the limit as $m\rightarrow\infty$ in (C.2) to find:
@@
$=\lim_{m\rightarrow\infty}\min_{\mathsf{P}\in\mathcal{P}}\mathsf{P}^{*}_{(x_{n_{m}},\pi_{n_{m}})}\cdot r^{(x_{n_{m}},\pi_{n_{m}})}.$
By Lemma C.4 we have $\mathsf{P}^{*}_{(x_{n_{m}},\pi_{n_{m}})}\rightarrow\mathsf{P}^{*}_{(x,\pi)}$ uniformly. Thus, $\lim\min=\min\lim$ implies,

### v2

@@ Appendix C Irreducible DR-MGs
for some positive distribution matrix $\mu$. Similarly, it implies $x(s)=x(s^{\prime})$, and thus $h^{\pi^{\star}}_{\mathsf{P}^{\star}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}=ce$. This further implies that all inequalities in (42) are equations, thus
$h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)=\max_{a}\left\{r(s,a)-g^{\pi^{\star}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})\right\},$ (45)
hence $(h^{\pi^{\star}}_{\mathsf{P}^{\star}},g^{\pi^{\star}}_{\mathsf{P}^{\star}})$ is a solution to (8), completing the proof of the first part. The remaining proof follows in a similar way from (wang2023model). ∎
@@
If for any agent $i$ and any joint policy $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, convex, and $\text{BR}_{i}(\cdot)$ is upper hemi-continuous, then BR also satisfies the conditions in Lemma 4.2.
@@
(3). Consider two convergent policy sequences $\{\pi^{n}\}\to\pi$ and $\{\mu^{n}\}\to\mu$, with $\mu^{n}\in\text{BR}(\pi^{n})$, we need to show that $\mu\in\text{BR}(\pi)$. Note that $\pi^{n}\to\pi$ implies that $\pi^{n}_{i}\to\pi_{i}$ for any $i$, and similarly $\mu^{n}\to\mu$. Since $\mu_{i}^{n}\in\text{BR}_{i}(\pi_{-i}^{n})$, by the hemi-continuity of $\text{BR}_{i}$, we have $\mu_{i}\in\text{BR}_{i}(\pi_{-i})$. Moreover note that
@@ C.1 Convexity
Now from Theorem 4.1, $g=g^{\pi}=g^{\star}_{\mathcal{P}}$, thus it holds that
@@
Define $\Pi^{\star}\triangleq\{\pi\ |\ \text{optimal policies of }g^{\pi}_{\mathcal{P}}\}$. Then $\Pi^{\star}$ is convex.
@@
Let $\pi_{1},\pi_{2}\in\Pi^{\star}$ and define $\tilde{\pi}=\alpha\pi_{1}+(1-\alpha)\pi_{2}$ for some $\alpha\in[0,1].$ Then by Theorem 4.1,
@@
Consequently, $(g,h)$ is also a solution to the robust Bellman equation w.r.t. $\tilde{\pi}$. This implies that $g^{\tilde{\pi}}=g=g^{\star}$, and thus that $\tilde{\pi}$ is also optimal.
@@ C.2 Hemi-continuity
Consider a policy sequence $\{\pi_{n}\}\rightarrow\pi$, and the sequence of induced transition matrices $\{\mathsf{P}_{\pi_{n}}\}$. It holds that $\mathsf{P}_{\pi_{n}}^{\star}\rightarrow\mathsf{P}_{\pi}^{\star}$ uniformly on $\mathcal{P}$.
@@
See Prop 8.4.6 in (puterman1994markov) for the proof. ∎
@@
Theorem C.2 (Hemi-continuity).
Let $\pi_{n}\rightarrow\pi$ and $x_{n}\rightarrow x$, where $\pi_{n}:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{-i})$ and $x_{n}:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{i})$. If $x_{n}\in\text{BR}_{i}(\pi_{n}),$ then $x\in\text{BR}_{i}(\pi).$
@@
Moreover, since $h_{n}$ is some relative function (wang2023robust), $h_{n}||_{\infty}\leq H.$ Then, there exists some subsequence $\{n_{m}\},$ s.t. $h_{n_{m}}\rightarrow h$ for some $h$, and $g^{(x_{n},\pi_{n})}_{\mathcal{P}}\rightarrow g$ for some $g$. Then, we can take the limit as $m\rightarrow\infty$ in (C.2) to find:
@@
$=\lim_{m\rightarrow\infty}\min_{\mathsf{P}\in\mathcal{P}}\mathsf{P}^{\star}_{(x_{n_{m}},\pi_{n_{m}})}\cdot r^{(x_{n_{m}},\pi_{n_{m}})}.$
By Lemma C.4 we have $\mathsf{P}^{\star}_{(x_{n_{m}},\pi_{n_{m}})}\rightarrow\mathsf{P}^{\star}_{(x,\pi)}$ uniformly. Thus, $\lim\min=\min\lim$ implies,